Supplemental Operations Report May 8th

(This is formatted for and is also being sent to the webteam, since we no longer have COMMS)

Supplemental Operations Report 8 May 2022

Name of person filing report: Shannon Rupert

Reason for Report: Routine

Non-nominal systems: Nothing to report

Action taken for non-nominal systems: nothing to report

Generator: Being manually run at night through Friday night, then offline and we are running on solar only

ScienceDome Dual Split: Off

Solar—Charge controllers off at night until Friday

Solar— Nominal

Notes on power system: nothing to report

Propane Reading, station tank – not noted %

Propane Reading, director tank— 80 %

Propane Reading, intern tank— 70 %

Propane Reading, generator—55 %

Ethanol Free Gasoline – 0 gallons.

Water (Outpost tank) – 500 gallons

Sojourner rover used: yes

Hours: not noted

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: (Honda, 350.1, 350.2, 300): 350.1 reinstalled as an additional sign for no trespassing at the entrance to the station

Reason for use: Hoping to stop trespassing

Oil Added? No.

ATV Fuel Used: 0 gallons

# Hours the ATVs were Used: 0 hours

Notes on ATVs: Nothing to report

HabCar used and why, where? Yes, moved around campus. The crew helped me reinstall the rear window and we used gorilla tape to secure it in place until it can be properly fixed.

CrewCar used and why, where? Yes, to town and to Grand Junction to drop off crew and return to the station. On Thursday I emptied the power steering compartment. The old power steering fluid was black. Once the compartment was empty, I refilled it with about 20 oz of clear power steering fluid. This should be done again at the beginning of the fall field season.

Luna used and why, where? No

Campus wide inspection, if action taken, what and why? Nothing to report

General notes and comments: I removed the tarps from the tunnel between the RAM and Hab and will continue to remove the remainder this week.

Summary of general operations: Nothing to report

Summary of internet: I took Starlink from the Hab to my house before the crew left. I’ll keep it here until the Work Party.

Summary of suits and radios: Spacesuits and radios are no longer being charged.

Summary of Hab operations: There is only about 50 gallons of water in the static tank and the pump is exposed. I will add more water as soon as I can. There was evidence of rodents in the rear airlock where they sampled the trash.

Summary of Outpost operations: Thanks to Isai and Bennie from Crew 265, the fence was completed on Saturday. They also removed the remainder of the fence and took the rocks to fit the barrier for the replaced conduit. My entire weekend of planned work was completed in about an hour! It was awesome! Update on Sunday: The old mattress next to the trash trailer was blown onto the fence and destroyed the new fence. God forbid I ever actually get ahead here!

I replaced and reinstalled the new pump in the water tank and it’s working perfectly.

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: I’m storing the tarps in here until I can get some help folding them.

Summary of any observatory issues: I still need to go into the robotic observatory and take the cameras off the telescopes. This work has been rescheduled for Wednesday, God willing and the creeks don’t rise.

Summary of health and safety issues: Nothing to report

Operations Report 2021-2022.docx
Operations Report 2021-2022.docx

Copyright © The Mars Society. All rights reserved. | Main Site